# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

### v1

@@ III-A The set of all error minimizing measurements and an arbitrary construction
and the equality is obtained iff measurement operators $\{\Lambda_{\rho},\Lambda_{\sigma}\}$ satisfy the condition:
$\begin{cases}\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\hfill\qquad\underline{\mathcal{C}1}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\hfill\qquad\underline{\mathcal{C}2}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{otherwise},\hfill\qquad\underline{\mathcal{C}3}\end{cases}$ (11)
where $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$.
@@
The proof is given in Appendix B. ∎
Recall from (6) that, minimum of $e(\Lambda)$ can be obtained by first finding $\sup_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$. It turns out that
@@
The first case in (11) is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies the condition that $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}>\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. In this case $\Lambda_{\rho}$ is taken such that $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$ and $\max_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}=\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma)}=R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)$. Similarly, second case is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies the condition that
$\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}<\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. The third case is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies when $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}=\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$, either can be chosen. We have to deal with the cases separately and hence need to represent them compactly. We represent them as $\mathcal{C}1,\mathcal{C}2,$ and $\mathcal{C}3$ respectively for the remaining of this section, as mentioned in (11).
@@
$\mathrm{T}^{\max}$ denotes the projection operator onto the subspace where $\rho$ is largest as comparison to $\sigma$, in the sense that any vector $|\psi\rangle$ in this subspace, we get $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\max}(\rho,\sigma)$, which is the highest possible value it can have. Similarly for the subspace corresponding to the projection operator $\mathrm{T}^{\min}$, this ratio $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\min}(\rho,\sigma)$, which is also the minimum possible value it can achieve. In classical scenario, on taking density operators corresponding to probabilities to be discriminated, $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ would correspond to choosing the realization of random variable for which ratio of the two probabilities are maximum and minimum.
@@
Now, $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\max}$. Similarly, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\min}$. So, Theorem III.1 states that, by restricting the subspace where eigenvector of the operators $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ lie, one can obtain measurements that achieve minimum postselected symmetric error.
Corollary 1.
If $p,\rho,\sigma$ satisfy $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, then for any $\Lambda$ such that $e(\Lambda)=e_{s}(\rho,\sigma,p)$, one of the following two holds:
  1. 1.
    The measurement never detects $\sigma$ that is $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$. The measurement outcome is either $\rho$ or third outcome i.e. rejecting both the hypotheses (see Figure 3(a)).
  2. 2.
    The measurement never detects $\rho$ that is $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0$. The measurement outcome is either $\sigma$ or third outcome i.e. rejecting both the hypotheses (see Figure 3(b)).
So, for $p,\rho,\sigma$ such that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, any measurement that minimizes postselected symmetric error, either never makes a decision in favour of $\rho$ or never makes a decision in favour of $\sigma$. Fig. 3 depicts the two possible measurements in this case.
Proof.
From the Theorem III.1, when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ we get
So, for any error-minimizing measurement, one of the two cases must hold. ∎
Figure 3: The figure shows the two possible outcomes when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ for any error minimizing measurement where (a) $\sigma$ never declared, (b) $\rho$ never declared.
Remark 3.
The statement in the Corollary 1 says that the measurements that achieve minimum postselected error, are not detecting one of the states. The observation puts a very fundamental question,“Does a lower postselected error really indicate better hypothesis testing, when prior probability is known?" The observation suggests a negative answer. So, the next question is, “What is a better metric to assess the quality of hypothesis testing, when prior probability is known?" We won’t address it in this work and leave it as an open question.
Corollary 2.
A measurement with non-zero probabilities of detecting both $\rho$ and $\sigma$ is possible only if the prior probabilities $p_{\rho}$ and $p_{\sigma}$ are equal to
Proof.
From (11), we know that such a measurement is possible when
@@
Now, putting the condition that $p_{\rho}+p_{\sigma}=1$, we obtain the desired result. ∎
@@
Figure 4: An illustration showing possible outcomes for different values of prior probability $p_{\rho}$ and $p_{\sigma}$.
From Theorem III.1, we now get three scenarios:
  • $p_{\rho}>p_{\rho}^{*}$ (i.e. $p_{\sigma}<p_{\sigma}^{*}$): Error-minimizing measurement never detects $\sigma$. It either detect $\rho$ or rejects both hypotheses. See Fig. 4(a).
  • $p_{\rho}<p_{\rho}^{*}$ (i.e. $p_{\sigma}>p_{\sigma}^{*}$): Error-minimizing measurement never detects $\rho$. It either detect $\sigma$ or rejects both hypotheses. See Fig. 4(b).
  • $p_{\rho}=p_{\rho}^{*}$ (i.e. $p_{\sigma}=p_{\sigma}^{*}$): There are error-minimizing measurements detecting either/both the hypotheses. See Fig. 4(c).
The observation says that if the prior probability of state being $\rho$ is high, $\sigma$ is never detected under the error minimizing measurement. Similarly, if the prior probability of state being $\sigma$ is high, $\rho$ is never detected. This states that an error minimizing measurement will detect the high probability state or declare nothing. Also, there exists a certain value of the pair $(p_{\rho},p_{\sigma})$ such that an error minimizing measurement can detect both states.
Now, we give two examples. In the first classical example, we show how the previous theorem can be used to find the error minimizing measurements. We also observe that there are measurements, despite having the minimum postselected symmetric error, rejecting both the hypotheses with probability close to $1$. This shows a need to find measurements that maximize acceptance.
In the first example, we take Hilbert’s space with basis given by $\{|0\rangle,|1\rangle,|2\rangle\}$. In this example, we get an intuitive idea when both state $\rho$ and $\sigma$, being a mixture of 3 orthogonal pure states $\{|0\rangle,|1\rangle,|2\rangle\}$. The first state $|0\rangle$ has higher probability in the state $\rho$ as compared to in the state $\sigma$, second orthogonal state $|1\rangle$ has higher probability in the state $\sigma$ as compared to state $\rho$ and third orthogonal state $|2\rangle$ having equal probability in both the states $\rho$ and $\sigma$. In relative terms, observe that, $|0\rangle$ is relatively prominent in the state $\rho$, $|1\rangle$ being prominent in the state $\sigma$, while $|2\rangle$ having equal probability of both. We obtain that the measurements that minimizes error, declares the unknown state as $\rho$ by measuring $|0\rangle\langle 0|$ and $\sigma$ is never detected. We will observe that, for the error-minimizing measurements, the value of acceptances for states $\rho$ and $\sigma$ vary depending on the choice of measurement. For an error-minimizing measurement, the acceptance can be very small, thus not declaring any state with probability close to 1. We suggest that, while designing an error-minimizing measurement, maximizing acceptance should also be considered.
@@
Let $\{|0\rangle,|1\rangle,|2\rangle\}$ be the basis of Hilbert’s space. Take $p_{\rho}=1/2,p_{\sigma}=1/2$,
$\rho=\frac{\mu}{2}|0\rangle\langle 0|+\frac{\mu}{2}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|\text{ and }\sigma=\frac{\mu}{4}|0\rangle\langle 0|+\frac{3\mu}{4}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|.$
Note that, here $\Pi_{\rho}=\Pi_{\sigma}=|0\rangle\langle 0|+|1\rangle\langle 1|+|2\rangle\langle 2|=\mathrm{I}$, so $\mathrm{I}-\Pi_{\sigma}=0$. Now,
@@
Thus minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=1/3$. Note that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)>R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, hence, it corresponds to the case $\mathcal{C}1$. We get
@@
So $\Lambda_{\rho}=c\frac{|0\rangle\langle 0|}{\mu/4}=\frac{4c}{\mu}|0\rangle\langle 0|$, $\Lambda_{\sigma}=0$. So, the set of POVM characterizing the measurements that achieve minimum postselected symmetric error is given by
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}=\frac{4c}{\mu}|0\rangle\langle 0|,\Lambda_{\sigma}=0,c\leq\frac{\mu}{4}\right\}$
=$=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:0<c\leq 1,\Lambda_{\rho}=c|0\rangle\langle 0|,\Lambda_{\sigma}=0\right\}.$ (19)
@@
Note that, with this POVM, a conclusive decision is being made for only $c\mu/2$ and $c\mu/4$ fraction of cases when the unknown state is $\rho$ and $\sigma$ respectively. If $\mu$ is small, performance of postselected test would be the same i.e. $e_{s}(\rho,\sigma,p)=1/3$, however we will be declaring inconclusive results in most of the cases, so the measurement would not be really useful. We need some metric to characterize the usefulness of postselected measurement. Second problem is choice of the constant term $c$. We use the metric acceptance to characterize the usefulness.
Now, we give second example. This quantum example is focused towards determining constraints and illustrating various claims made in this subsection about error minimizing measurements.

### v2

@@ 3. $\bullet$ Case $\underline{\mathcal{C}\mathrm{3}}$:
and the equality is obtained iff the measurement operators $\{\Lambda_{\rho},\Lambda_{\sigma}\}$ satisfy the condition:
$\begin{cases}\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0,&\qquad\text{for Case $\underline{\mathcal{C}1}$,}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\qquad\qquad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\qquad\text{for Case $\underline{\mathcal{C}2}$,}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\qquad\text{for Case $\underline{\mathcal{C}3}$,}\end{cases}$ (11)
where $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$, and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$.
@@
The proof is given in Appendix A. ∎
Recall from (6) that the minimum value of $e(\Lambda)$ requires finding $\sup_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$. We substitute $\max$ in the place of $\sup$, as the maximum is shown to be achievable in [33]. We can express the maximum as
@@
The first case in (11) corresponds to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies the condition that $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}>\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. In this case, $\Lambda_{\rho}$ is taken such that $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$ and $\max_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}=\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma)}=R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)$. Similarly, second case is corresponds to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, which signifies the condition that
$\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}<\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. The third case corresponds to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies when $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}=\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$, either can be chosen. We will have to deal with the cases separately.
@@
$\mathrm{T}^{\max}$ denotes the projection operator onto the subspace where $\rho$ is largest in comparison to $\sigma$, in the sense that any vector $|\psi\rangle$ in this subspace, we get $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\max}(\rho,\sigma)$, which is the highest possible value it can have. Similarly, for the subspace corresponding to the projection operator $\mathrm{T}^{\min}$, this ratio $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\min}(\rho,\sigma)$, which is also the minimum possible value it can achieve. In the classical scenario, i.e., while discriminating a pair of probability distributions, corresponding to some random variable, on taking density operators corresponding to probabilities to be discriminated, $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ would respectively correspond to choosing the realization of the random variable for which the ratio of the two probabilities are maximum and minimum.
@@
Now, $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\max}$. Similarly, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\min}$. So, Theorem 1 states that, by restricting the subspace where eigenvector of the operators $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ lie, one can obtain measurements that achieve the minimum postselected symmetric error.
Now note that the condition $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ in $\mathcal{C}3$ is equivalent to
@@
(15) along with $p_{\rho}+p_{\sigma}=1$ gives $p_{\rho}=p_{\rho}^{*}$ and $p_{\sigma}=p_{\sigma}^{*}$ where
Further, we can show that the condition $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)>R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ is equivalent to $p_{\rho}>p_{\rho}^{*}$ (i.e., $p_{\sigma}<p_{\sigma}^{*}$). Hence, we can write the three cases in Theorem 1 as
  • $\underline{\mathcal{C}1}:p_{\rho}>p_{\rho}^{*}$ (i.e., $p_{\sigma}<p_{\sigma}^{*}$),
  • $\underline{\mathcal{C}2}:p_{\rho}<p_{\rho}^{*}$ (i.e., $p_{\sigma}>p_{\sigma}^{*}$),
  • $\underline{\mathcal{C}3}:p_{\rho}=p_{\rho}^{*}$ (i.e., $p_{\sigma}=p_{\sigma}^{*}$).
Corollary 1.
If $p,\rho,\sigma$ satisfy $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, or equivalently $p_{\rho}\neq p^{*}_{\rho}$, then for any $\Lambda$ such that $e(\Lambda)=e_{s}(\rho,\sigma,p)$, one of the following two holds:
  1. 1.
    The measurement never detects $\sigma$ that is $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$. The measurement outcome is either $\rho$ or the third outcome, i.e., rejecting the test(see Fig. 4 (a)).
  2. 2.
    The measurement never detects $\rho$ that is $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0$. The measurement outcome is either $\sigma$ or the third outcome, i.e., rejecting the test (see Fig. 4 (b)).
So, for $p,\rho,\sigma$ such that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, any measurement that minimizes postselected symmetric error, either never decides in favor of $\rho$ or never makes a decision in favor of $\sigma$. Fig. 4 (a) and Fig. 4 (b) depict the two possible measurements in this case.
Proof.
From the Theorem 1, when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ we get
So, one of the two cases must hold for any error-minimizing measurement. ∎
Remark 3.
The statement in the Corollary 1 says that a measurement that achieves the minimum postselected error can only detect one of the state. The observation puts a very fundamental question, “Does a lower postselected error indicate better hypothesis testing when prior probability is known?" The observation suggests a negative answer. So, the next question is, “What is a better metric to assess the quality of hypothesis testing when prior probability is known?" We will not address it in this work and leave it as an open question.
@@
Figure 4: An illustration showing possible outcomes for different prior probability values of $p_{\rho}$ and $p_{\sigma}$. The figure shows the possible outcomes when $p_{\rho}\neq p_{\rho}^{*}$ (equivalently $p_{\sigma}\neq p_{\sigma}^{*}$) for any error minimizing measurement where (a) $\sigma$ never declared, (b) $\rho$ never declared. Further, (c) shows that if $p_{\rho}=p_{\rho}^{*}$ (equivalently $p_{\sigma}=p_{\sigma}^{*}$), all the outcomes are possible.
Corollary 2.
A measurement with non-zero probabilities of detecting both $\rho$ and $\sigma$ is possible only if the prior probabilities $p_{\rho}=p_{\rho}^{*}$ and $p_{\sigma}=p_{\sigma}^{*}$. Further
  • $\underline{\mathcal{C}1}:p_{\rho}>p_{\rho}^{*}$ (i.e., $p_{\sigma}<p_{\sigma}^{*}$): Error-minimizing measurement never detects $\sigma$. It either detects $\rho$ or rejects the test. See Fig. 4(a).
  • $\underline{\mathcal{C}2}:p_{\rho}<p_{\rho}^{*}$ (i.e., $p_{\sigma}>p_{\sigma}^{*}$): Error-minimizing measurement never detects $\rho$. It either detects $\sigma$ or rejects the test. See Fig. 4(b).
  • $\underline{\mathcal{C}3}:p_{\rho}=p_{\rho}^{*}$ (i.e., $p_{\sigma}=p_{\sigma}^{*}$): There are error-minimizing measurements detecting either/both the hypotheses. See Fig. 4(c).
The observation says that if the prior probability of a state being $\rho$ is high, $\sigma$ is never detected for any error-minimizing measurement. Similarly, if the prior probability of the state being $\sigma$ is high, $\rho$ is never detected. So, an error-minimizing measurement will detect the high probability state or declare nothing. Also, a particular value of the pair $(p_{\rho},p_{\sigma})$ exists such that an error-minimizing measurement can detect both states.
For better exposition, we now give two examples. The first example shows how Theorem 1 can be used to find the error-minimizing measurements. We also observe that there are measurements, despite having the minimum postselected symmetric error, rejecting both hypotheses with a probability close to $1$. The observation shows that all error-minimizing measurements are not useful and there is a need to find those that also maximize acceptance.
Here, we take a Hilbert’s space with it’s basis $\{|0\rangle,|1\rangle,|2\rangle\}$. We consider both $\rho$ and $\sigma$ to be a mixture of 3 orthogonal pure states $\{|0\rangle,|1\rangle,|2\rangle\}$. $\rho$ has a higher probability to be in $|0\rangle$ than the probability of $\sigma$ to be in $|0\rangle$. Further, $\rho$ has a smaller probability to be in $|1\rangle$ than $\sigma$. $|2\rangle$ constitutes $\rho$ and $\sigma$ with equal probability. In relative terms, observe that $|0\rangle$ is relatively prominent in the state $\rho$, $|1\rangle$ being prominent in the state $\sigma$, while $|2\rangle$ has an equal probability of both.
@@
Let $\{|0\rangle,|1\rangle,|2\rangle\}$ be the basis of a Hilbert’s space. Let us consider that two states
$\rho=\frac{\mu}{2}|0\rangle\langle 0|+\frac{\mu}{2}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|\text{ and }\sigma=\frac{\mu}{4}|0\rangle\langle 0|+\frac{3\mu}{4}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|$
with prior probabilities $p_{\rho}=1/2,p_{\sigma}=1/2$. Note that, here $\Pi_{\rho}=\Pi_{\sigma}=|0\rangle\langle 0|+|1\rangle\langle 1|+|2\rangle\langle 2|=\mathrm{I}$, so $\mathrm{I}-\Pi_{\sigma}=0$. Now,
@@
Thus, the minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=1/3$. Note that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)>R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, so $\mathcal{C}1$ is applicable here. Hence, from Theorem 1, we have
@@
So $\Lambda_{\rho}=c\frac{|0\rangle\langle 0|}{\mu/4}=\frac{4c}{\mu}|0\rangle\langle 0|$, $\Lambda_{\sigma}=0$. So, the set of POVM characterizing the measurements that achieve the minimum postselected symmetric error is given by
$\mathcal{E}_{s}(\rho,\sigma,p)$ $=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}=\frac{4c}{\mu}|0\rangle\langle 0|,\Lambda_{\sigma}=0,c\leq\frac{\mu}{4}\right\}$
$=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:0<c\leq 1,\Lambda_{\rho}=c|0\rangle\langle 0|,\Lambda_{\sigma}=0\right\}.$ (19)
@@
Note that, with this POVM, a conclusive decision is being made for only $c\mu/2$ and $c\mu/4$ fraction of cases when the unknown state is $\rho$ and $\sigma$ respectively. If $c$ is small, the performance of the postselected test would be the same, i.e., $e_{s}(\rho,\sigma,p)=1/3$; however, we will be declaring inconclusive results most of the time, so the measurement would not be useful. We need some metrics to characterize the usefulness of postselected measurement. We use the metric acceptance to characterize the usefulness.
In the above example, we saw that all error-minimizing measurements declare the unknown state as $\rho$ if the outcome is $|0\rangle\langle 0|$, otherwise they reject the test. Therefore, $\sigma$ is never detected. We observe that, for the error-minimizing measurements, the value of acceptances for states $\rho$ and $\sigma$ vary depending on the choice of measurement, in particular, on the choice of the parameter $c$. For an error-minimizing measurement, the acceptance can be very small, which means that the measurement makes decision with a probability close to 0. The example suggests that maximizing acceptance should also be considered while designing an error-minimizing measurement. The second example is designed to explain the derived constraint and illustrates various observations made in this subsection about error-minimizing measurements.